Complete the following balance sheet for the Range Company using the following information:

Debt to Assets = 60%
Quick Ratio = 1.1
Asset Turnover = 5x
Capital Asset Turnover = 12.037x
Current Ratio = 2
Average Collection Period = 17.0708 days

Assume all sales are on credit.

Cash                                        _______          Current Liabilities                  _______

Receivables                            _______          Bonds Payable                        _______

Inventory                                _______          Total Liabilities                      _______

Total Current Assets             _______          Net Worth                               _______           

Plant and Equipment              _______

Total Assets                          $325,000            Total Liab and NW              _______
Answer #1

v

Total assets: 325000
Debt to assets: 60%
Debts = 325000*60% = 195000
Net worth =Total assets - Debts = 325000-195000 = 130000
Assets turnover = 5x
Sales revenue = 5*325000 = 1625000
Capital Assets Turnover = 12.037x
Capital assets (Fixed assets)= 1625000/12.037 =135000
Current assets = Total assets - Fixed assets = 325000-135000 = 190000
Current ratio = 2
Current ratio = Current assets / Current liabilities
2 = 190000 / Current liabilities
Currrent liabilities = 95000
Total debts = Current liabilities + Bonds payable
Bonds payable = 195000-95000 = 100000
Quick ratio = 1.1
Quick ratio = Liquid assets / Curren liabilities
1.1 = Liquid assets / 95000
Liquid assets = 104500
Inventory = Current assets -Liquid assets
Inventory = 190000-104500 = 85500
Average collection period = 17.0708 days
Average collection period = 365 *Accounts receivable / Net sales
17.0708 = 365 / 1625000 * Accounts receivable
Accounts receivable = 76000
Current assets = Cash+ Accounts receivable + Inventory
190000 =Cash + 76000+85500
Cash = 28500
Balance Sheet:
Cash 28500 Current liabilities 95000
Receivables 76000 Bonds payable 100000
Inventory 85500 Total liabilities 195000
Total Current assets 190000
Plant and equipment 135000 Net worth 130000
Total assets 325000 Total liab and NW 325000